OpenStudy (anonymous):

The jogging area in a club is in the shape of the trapezium. If the trapezium is based on a rectangular coordinate system, what is the square root of its area if the imaginary vertices have the coordinates (2,-2),(2,3),(20,2),(20,-2)? 7.5, 9, 11, 13

OpenStudy (anonymous):

area = 0.5 (5+4) * 18 = 4.5 * 18 = 81 sqrt 81 = 9 9 is your answer

OpenStudy (anonymous):